OpenStudy (anonymous):

log9(x+6)-log9x=log9 2

OpenStudy (anonymous):

9 is base I'm just asking whether this is even solvable

OpenStudy (zzr0ck3r):

hint \(\log_9(x+6)-\log_9(x)=\log_9(2)\\\log_9(\frac{x+6}{x})=\log_9(2)\\\frac{x+6}{x}=2\)

OpenStudy (zzr0ck3r):

can you solve that?

OpenStudy (anonymous):

that's what i ended up with

OpenStudy (zzr0ck3r):

good work

OpenStudy (anonymous):

well then what do you put as final answer just that?

OpenStudy (zzr0ck3r):

you can always check to see if your right here http://www.wolframalpha.com/input/?i=log9%28x%2B6%29-log9%28x%29%3Dlog9%282%29

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

solve for x

OpenStudy (zzr0ck3r):

\(\dfrac{x+6}{x}=2\)

OpenStudy (zzr0ck3r):

multiply both sides by \(x\) \(x+6=2x\)

OpenStudy (anonymous):

u end up with 6=2?

OpenStudy (anonymous):

oooh

OpenStudy (zzr0ck3r):

no of course not, 6 is not equal to 2

OpenStudy (anonymous):

ima dumbass

OpenStudy (zzr0ck3r):

if you are having problems solving this, I would go back and review old material. You should not be working on logs if you cant solve this IMO

OpenStudy (zzr0ck3r):

\(x+6=2x \implies x=6\)

OpenStudy (anonymous):

lol thanks for the help